OpenStudy (anonymous):

Choose the correct simplification of 5^4 over 5^2

OpenStudy (diyadiya):

\[\Large \frac{x^m}{x^n}= x^{m-n}\] So \[\Large \frac{5^4}{5^2}= 5^{4-2}= ?\]